The measure of angle b, the supplement of angle a, is four times the measure of angle c, the complement of angle a.
Answer by fractalier(2101)

We can deduce these algebraic facts from yur statements...
b + a = 180
b = 4c
c + a = 90
If we substitute 4c for b in the first equation, we get
4c + a = 180
c + a = 90
Now subtract equations and get
3c = 90
c = 30
b = 120
a = 60
and all checks out.